OpenStudy (anonymous):

Find the dimensions of a right-circular cylinder that is open on the top and closed on the bottom, so that the can holds 1 liter and uses the least amount of material?

OpenStudy (anonymous):

seems like this has lack of info volume of cylinder v=pi *h*r^2 , h=height,r=radius or v=pi*h*d^2 /4

OpenStudy (anonymous):

Nope, this is it. It is from Calculus, hence involves derivation.

OpenStudy (anonymous):

for the rectangle, circumference of circle C=2pi*r, its area(of rectangle) A1=C*h=2pi*rh while area of the bottom circle A2=pir^2 now the total area A=A1+A2 A=2pi*rh+pi*r^2, derivative with respect to r dA/dr=2pi*h+2pi*r=0 to (get the minimum) solve for r r=-h to get the min material we need to use r=-h now volume V=pi*hr^2 V=pi*h(-h)^2 V=pi*h^3 iliter=pi*h^3 height h=cube root(1liter/pi) h=0.683, also radius r=0.683

OpenStudy (dumbcow):

Goal is to minimize surface area, given volume of 1 liter. 1 liter = 1000 mL = 1000 cubic cm So i will assume they want dimensions in terms of cm. \[\pi*r^{2}*h = 1000\] \[SA = \pi*r^{2} + 2\pi*r*h\] From volume equation solve for h \[h = \frac{1000}{\pi*r^{2}}\] Substitute into surface area equation \[SA = \pi*r^{2} + 2\pi*r*(\frac{1000}{\pi*r^{2}}) = \pi*r^{2} + \frac{2000}{r}\] differentiate \[dSA/dr = 2*\pi*r - \frac{2000}{r^{2}}\] set equal to zero and solve for r \[2*\pi*r - \frac{2000}{r^{2}} = 0\] \[\rightarrow 2*\pi*r^{3} = 2000\] \[\rightarrow r^{3} = \frac{1000}{\pi}\] \[r \approx 6.83\] Substitute back in to solve for h \[h = \frac{1000}{\pi*(6.83)^{2}} = 6.83\]

OpenStudy (anonymous):

yup we got the same answer..lol

OpenStudy (dumbcow):

yep checks out...not sure on units though what units would yours be in?

OpenStudy (anonymous):

Thank you so much guys! xD

OpenStudy (dumbcow):

your welcome...good luck on the exam

OpenStudy (anonymous):

international units or english measures will work ,,,lol

OpenStudy (anonymous):

Thanks, once again! =]

OpenStudy (anonymous):

Hey 'dumbcow' (sorry I don't know your name lol) can you please look at this one: http://openstudy.com/groups/mathematics/updates/4e376adf0b8bf47d065ed79d